OpenStudy (clueless):

A circle is centered at the point (-3, 2) and passes through the point (1, 5). The radius of the circle is ___ units. The point (-7, ___) lies on this circle.

jimthompson5910 (jim_thompson5910):

Use the distance formula to find the radius (x1,y1) = (-3,2) x1 = -3 y1 = 2 (x2,y2) = (1,5) x2 = 1 y2 = 5 \[\Large d = \sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\] \[\Large d = \sqrt{(-3-1)^2+(2-5)^2}\] I'll let you finish

OpenStudy (clueless):

so the radius is -8? or -64?

jimthompson5910 (jim_thompson5910):

the radius will be some positive number

jimthompson5910 (jim_thompson5910):

\[\Large d = \sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\] \[\Large d = \sqrt{(-3-1)^2+(2-5)^2}\] \[\Large d = \sqrt{(-4)^2+(-3)^2}\] \[\Large d = \sqrt{16+9}\] \[\Large d = ???\]

OpenStudy (clueless):

5?

jimthompson5910 (jim_thompson5910):

d = 5, yes

jimthompson5910 (jim_thompson5910):

the distance is 5, so the radius is r = 5

OpenStudy (clueless):

then would the point lie on -7, -2?

jimthompson5910 (jim_thompson5910):

there are two points on the red circle such that the x coord is -7 http://prntscr.com/bax2wr

jimthompson5910 (jim_thompson5910):

`then would the point lie on -7, -2?` no, not correct

OpenStudy (clueless):

so it's either -7, -1 or -7, 5?

jimthompson5910 (jim_thompson5910):

use parenthesis whenever talking about ordered pairs yes (-7,-1) or (-7,5)

OpenStudy (clueless):

Yay! I think I'm finally starting to understand this stuff more, thank you for helping me out so much.

jimthompson5910 (jim_thompson5910):

you're welcome